# Controllable Prompt Tuning For

### v1

@@ 1 Introduction
Spurious correlation or shortcut learning arises when a classifier relies on non-predictive features that coincidentally correlate with class labels among training samples. For example, a majority of waterbirds are often found near water, while landbirds mostly appear near land. Even when being provided with a small number of minority group examples, such as waterbirds on land, or landbirds on water, model predictions are often still dominated by spurious features, in this case, the background, ignoring the salient feature, the foreground [68].
A variety of approaches aim to improve worst group performance, by oversampling high-loss samples [53, 90], undersampling the majority group [69] or re-training a last layer on group-balanced validation data [42]. A particularly popular and effective baseline, Group Distributionally Robust Optimization (GroupDRO) [68, 69], directly minimizes an estimated upper bound on worst group loss. However, all these procedures generally neglect knowledge transfer between groups. Moreover, high-error samples can potentially include noisy-label data that could hurt the model’s predictive ability if we try to learn them exhaustively [61]. Besides, DRO-based methods, in particular, are susceptible to overfitting, wherein their test performances can decline to the same as ERM with sufficiently long training [31, 64, 88].
Figure 1: Accuracy curves on Waterbirds for four groups during training. Vertical lines indicate early stopping epochs when models perform best on the validation set.
@@
Moreover, we aim to control the group-wise learning process by ensuring that the magnitude of loss for different groups is inversely proportional to a predefined vector $\mathbf{c}$. Varying this hyperparameter $\mathbf{c}$ adjusts the trade-off between the worst group and the average performance, yielding a more flexible method than prior work that only focuses on maximizing the worst group accuracy. However, computation can grow linearly with the number of controlling vectors $\mathbf{c}$. To overcome this challenge, we propose decoupling our optimization framework with parameter-efficient fine-tuning (PEFT) techniques that considerably reduce the number of trainable parameters. With the model compression ability powered by prompt-tuning, our proposed model can scale well for a large number of controlling vectors with a small parameter increase, e.g., $0.01\%$ and $0.003\%$ per value of $\mathbf{c}$ for Vision Transformer [23] and CLIP [66], respectively.
@@
  • Drawing inspiration from the principles of multi-objective optimization theory, we introduce a novel balancing mechanism to address the challenges of group distributional robustness. Our method not only considers the worst group but also leverages the gradient information from all groups to determine an ultimate updating direction that benefits all of them.
  • While the above rigid balancing procedure works well in different scenarios, we extend our method by introducing a controlling vector that allows us to dynamically adjust the priority across groups. Additionally, we integrate prompt tuning techniques to enhance the scalability of our method as the complexity and number of controlling vectors increase and to make it applicable for de-biasing large transformers-based models.
  • We conduct intensive experiments on many different benchmarks, where CPT consistently exhibits superior performance. Our method surpasses the current state-of-the-art baselines on Waterbirds and CelebA datasets while updating $0.4\%$ parameters. Moreover, it also outperforms recently proposed methods that aim to de-bias Vision Transformer and CLIP models with minimal training cost.
@@ 2 Related work
can be compromised for reasons like selection bias or spurious correlation across tasks. For selection bias on individual samples, prior works seek to reduce bias via practices like importance reweighting [36, 72, 16, 46], hard sample reweighting [53, 60] or distribution matching and discrepancy minimization [17, 7, 8, 63], and domain-adversarial algorithms [27, 58] across tasks in their feature representation space. For selection bias on groups or subpopulation, namely subpopulation shift or dataset imbalance, label propagation  [11, 8] or other consistency regularization [59, 84] are used to generalize the prediction to broader domains. In particular, GroupDRO [68, 69] or other relevant methods [44, 39] select descending direction more in favor of the worse performing/minority group(s) during training.
@@
Improving distributional robustness also requires mitigating the effect of environmental features. Prior work uses model ensemble [45] or model soups [81] to learn rich features and reduce the effect of spurious features. More practices include invariant feature representation learning [6, 14], domain generalization [44, 73, 71] or through data augmentation [83, 86].
@@
Transformer In the past few years, transformer [77] has emerged as a dominant architecture for a wide range of applications, exhibiting their profound capacity in modeling different modalities: languages [21, 55, 76], vision [23, 12], speech [67, 79], and time series [49, 30]. Due to its ability to process multiple modalities, transformer-based architectures were widely adopted for image understanding using different types of supervision. For example, Vision Transformer (ViT) [23] utilizes masked patches for pre-training and image-label supervision for fine-tuning. Notably, Contrastive Language-Image Pre-training (CLIP) [66], which was pre-trained on billion-scale image-text pairs, has demonstrated excellent zero-shot performance in image classification tasks. Despite obtaining impressive results on various downstream tasks, over-parameterization is still a crucial problem for transformer [26, 62] which potentially causes overfitting [48] and unintended biases [1, 80, 24, 89] more severely than small-scale architectures [69].
@@
Parameter-Efficient Fine-Tuning While conventional training updates the entire network, prior work has shown that it is possible to achieve performance comparable with full fine-tuning by instead updating a small set of parameters [51, 33, 34]. Early work in this line of research includes adapter [37, 13, 74], which adds a lightweight fully connected network between the layers of a frozen pre-trained model. Since then, there have been many other methods that update subsets of the network [32, 28, 87] or reparametrize the weights of the network using low-rank decompositions [2, 38, 25]. Another appealing PEFT strategy is prompt tuning, which prepends a few trainable tokens before the input sequence or hidden state to encode task-specific knowledge to the pre-trained model [47, 50, 94, 40].
@@ 3.1 Problem formulation
Following previous methods [69], we adopt two metrics: worst group accuracy, indicating the minimum test accuracy across all groups $g\in\mathcal{G}$, and average accuracy, which represents the weighted average test accuracy with the weights corresponding to the relative proportions of each group in the training data. It is worth noting that commonly used datasets for group robustness often exhibit skewed training set distributions, the weighted average score is thus dominated by the performance of the model on groups that include large numbers of training data. Therefore, we propose to evaluate the model performance by additionally reporting the mean (unweighted average) accuracy score across different groups.
@@ 3.3 Multi-objective Optimization
In general, an optimal solution for this objective vector function does not exist since each of the individual function $f_{i}(x)$ does not necessarily guarantee to have the same minimum solution. Alternatively, we expect to obtain a solution, from which we cannot improve any specific objective without hurting another. According to the above formulation, the solution [95] of the multi-objective minimization problem is formally defined as follows:
Definition 1.
@@
Gradient-based MOO Désidéri [20] shows that the descent direction can be found in the convex hull composed of gradients corresponding to different objectives $\{\boldsymbol{h}^{i}:=\nabla f_{i}\}_{i=1}^{m}$:
@@ 4 Proposed method
In GroupDRO, the updating formula solely focuses on the worst-performing group, which is susceptible to challenges when groups exhibit varying loss magnitude, noise, and transfer characteristics. For example, among the high-loss group, there might exist a few outliers that have large input shifts or even wrong labels, which can raise the training loss of the whole group and prevent the model from fitting other groups.
@@ 4.1 Balancing group distributional robustness
Figure 2: Overview of our method on the Waterbirds dataset. Our main objective is not only to improve model performance across groups by optimizing their loss functions $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$, but also to maximize the entropy over this loss distribution.
@@
However, MGDA is often biased toward objectives with smaller gradient magnitudes [54], and lacks controllable property (i.e. could not adjust the group that we want to focus more or less at each update), which can be utilized to balance the loss between groups. Hence, we introduce a new objective that steers the updating direction to obtain our desired balanced magnitude among groups. Assuming that the entropy $H(p)$ measures the diversity of a continuous distribution $p$, this behavior could be achieved via maximization of the entropy with respect to the distribution of loss functions:
$\mathcal{L}_{ent}(\theta)$ $=H\big{(}\text{softmax}\big{(}\vec{\mathcal{L}}(\theta))\big{)}=H\Big{(}\text{softmax}\big{(}\big{\{}\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]\big{\}}_{k=1}^{K}\big{)}\Big{)}$ (5)
Theorem 3.
@@
The proof of Theorem 3 is deterred to Appendix A. Theorem 3 provides us with the updating rule that increases the entropy term in Equation (5) the most. To balance the learning across groups, we seek the coefficient $w$ that maximizes the cosine similarity with $d_{ent}$ and does not conflict against the gradient vector of each group. Denotes $w^{\prime}_{i}=p_{i}\log(p_{i})-p_{i}\sum_{j=1}^{K}\log(p_{j})p_{j}$, this is equivalent to optimizing:
@@
While the balance between loss magnitude is not achieved, we strictly follow the updating direction that decreases training errors of high-loss groups but slightly relaxes constraints in the optimization problem (6) to allow bounded increases among groups that already have small loss magnitudes.
@@
In summary, the key difference between our proposed method and GroupDRO is that we solve a small $K$-dimensional optimization problem to find the updating direction that not only focuses on the worst group but also improves the performance in other groups and balances the loss magnitude between them at the same time. In practical settings, the group number $K$ is often a small number (e.g. $K=4$), thus could be efficiently solved using standard linear programming libraries [22, 5] without introducing significant computational overhead (Appendix B).
@@ 4.2 Controllable prompt tuning for balancing group distributional robustness
To achieve the controllable property, we apply the optimization procedure in section 4.1 on $\ell_{k}(\theta)={c}_{k}\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]$ where $\mathbf{c}\in\mathbb{R}^{K}$ is a pre-defined vector that is used to adjust the magnitude of group losses. Intuitively, the principle behind our proposed objective is to enforce the loss vector $\vec{\mathcal{L}}(\theta)$ to be inversely proportional to $\mathbf{c}$, from which we can adjust the group loss magnitude, and thus control the trade-off between them by varying $\mathbf{c}$. For example, we can put more weight on penalizing loss terms of some specific groups to accelerate the learning progress on those groups, or conversely slow down the learning to avoid overfitting on oversampled groups [19].
@@
Overview of our proposed method is depicted in Figure 2, given an input image belonging to one of a group $\{g^{1},g^{2},\dots,g^{K}\}$ and a set of classes, we solve the optimization problem (4.1) to find the reweighting coefficient $w$ that maximizes the entropy $\mathcal{L}_{ent}(\theta)$ while reasonably decreases group losses. While Figure 2 illustrates our proposed method on the CLIP model on the image classification task, it applies to other transformer-based models and task types. In our experiment, we exploit both image-end prompts and language-end prompts tuning on frozen ViT and CLIP backbones, respectively.
Prompt tuning for Vision Transformer: Regarding the prompt design for a ViT of $N$ layers, we use the set of continuous learnable tokens $\{\mathbf{P}_{n}\}_{n=0}^{N-1}$. Each $\mathbf{P}_{n}\in\mathbb{R}^{L\times D}$, where $D$, $L\in\mathbb{N}$ indicate the prompt length and latent space dimension, respectively. Subsequently, the input at the $(n+1)$-th layer follows the format of $\left[x_{n},\mathbf{P}_{n},\mathbf{E}_{n}\right]$, where $\mathbf{x}_{n},\mathbf{E}_{n}\in\mathbb{R}^{D}$ represent the [CLASS] token and image patch embeddings after $n$-th layer while [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operator. Thus, the number of trainable parameters for this image-end prompt set is $L\times N\times D$.
Prompt tuning for CLIP: Being trained on an enormous amount of image-text pairs, CLIP offers an expressive representation for both language and image inputs by using two separate encoders, one for each modality, as shown in Figure 2. Since the image encoder can be either ResNet [35], ConvNeXt [57] or ViT [23], we conduct prompt tuning on the transformer text encoder for consistency. In short, we employ a learnable context prompt of length $L$: $\mathbf{P}\in\mathbb{R}^{L\times D}$, to replace the hand-crafted prompt used by the original CLIP (i.e., “this is a photo of a "). Based on this design, the input for the text encoder of each class has the format of $\left[x,\mathbf{P}\right]$, where $x$ is the [CLASS] embedding of the corresponding class name and has the same dimension of $D$. These learnable tokens $\mathbf{P}$ are better at capturing domains-specific knowledge than those artificial prompt [92, 93] and can be learned to de-bias the original CLIP model while introducing a few trainable parameters ($N\times D$).

### v2

@@ 1 Introduction
Spurious correlation or shortcut learning arises when a classifier relies on non-predictive features that coincidentally correlate with class labels among training samples. For example, a majority of waterbirds are often found near water, while landbirds mostly appear near land. Even when being provided with a small number of minority group examples, such as waterbirds on land, or landbirds on water, model predictions are often still dominated by spurious features, in this case, the background, ignoring the salient feature, the foreground (Sagawa et al., 2019).
A variety of approaches aim to improve worst group performance, by oversampling high-loss samples (Liu et al., 2021b; Zhang et al., 2022), undersampling the majority group (Sagawa et al., 2020) or re-training a last layer on group-balanced validation data (Kirichenko et al., 2022). A particularly popular and effective baseline, Group Distributionally Robust Optimization (GroupDRO) (Sagawa et al., 2019, 2020), directly minimizes an estimated upper bound on worst group loss. However, all these procedures generally neglect knowledge transfer between groups. Moreover, high-error samples can potentially include noisy-label data that could hurt the model’s predictive ability if we try to learn them exhaustively (Oh et al., 2022). Besides, DRO-based methods in particular are susceptible to overfitting, wherein their test performances can decline to the same as ERM with sufficiently long training (Gulrajani & Lopez-Paz, 2020; Piratla et al., 2021; Zhai et al., 2022).
Figure 1: Accuracy curves on Waterbirds for four groups during training. Vertical lines indicate early stopping epochs as models obtain the best performance on the validation set.
@@
Moreover, we aim to control the group-wise learning process by ensuring that the magnitude of loss for different groups is inversely proportional to a predefined vector $\mathbf{c}$. Varying this hyperparameter $\mathbf{c}$ adjusts the trade-off between the worst group and the average performance, yielding a more flexible method than prior work that only focuses on maximizing the worst group accuracy. However, computation can grow linearly with the number of controlling vectors $\mathbf{c}$. To overcome this challenge, we propose decoupling our optimization framework with parameter-efficient fine-tuning (PEFT) techniques that considerably reduce the number of trainable parameters. With the model compression ability powered by prompt-tuning, our proposed model can scale well for a large number of controlling vectors with a small parameter increase, e.g., $0.01\%$ and $0.003\%$ per value of $\mathbf{c}$ for Vision Transformer (Dosovitskiy et al., 2020) and CLIP (Radford et al., 2021), respectively.
@@
$\bullet$ Drawing inspiration from the principles of multi-objective optimization theory, we introduce a novel balancing mechanism to address the challenges of group distributional robustness. Our method not only considers the worst group but also leverages the gradient information from all groups to determine an ultimate updating direction that benefits all of them.
$\bullet$ While the above rigid balancing procedure works well in different scenarios, we extend our method by introducing a controlling vector that allows us to dynamically adjust the priority across groups. Additionally, we integrate prompt tuning techniques to enhance the scalability of our method as the complexity and number of controlling vectors increase and to make it applicable for de-biasing large transformers-based models.
$\bullet$ We conduct intensive experiments on many different benchmarks, where CPT consistently exhibits superior performance. Our method surpasses the current state-of-the-art baselines on Waterbirds andimprove the performance CelebA datasets while updating $0.4\%$ parameters. Moreover, it also outperforms recent proposed methods that aims to de-bias Vision Transformer and CLIP models with minimal training cost.
@@ 2 Related work
can be compromised for reasons like selection bias or spurious correlation across tasks. For selection bias on individual samples, prior works seek to reduce bias via practices like importance reweighting (Heckman, 1979; Shimodaira, 2000; Cortes et al., 2010; Lei et al., 2021), hard sample reweighting (Liu et al., 2021b; Nam et al., 2020) or distribution matching and discrepancy minimization (Cortes et al., 2015; Ben-David et al., 2010; Berthelot et al., 2021), and domain-adversarial algorithms (Ganin et al., 2016; Long et al., 2018; Phan et al., 2023) across tasks in their feature representation space (Tran et al., 2023). For selection bias on groups or subpopulation, namely subpopulation shift or dataset imbalance, label propagation  (Cai et al., 2021; Berthelot et al., 2021) or other consistency regularization (Miyato et al., 2018; Yang et al., 2023a) are used to generalize the prediction to broader domains.
@@
Improving distributional robustness also requires mitigating the effect of environmental features. Prior work uses model ensemble (Kumar et al., 2022) or model soups (Wortsman et al., 2022) to learn rich features and reduce the effect of spurious features. More practices include invariant feature representation learning (Arjovsky et al., 2019; Chen et al., 2022b), domain generalization (Krueger et al., 2021; Sun et al., 2017; Shi et al., 2021) or through data augmentation (Xu et al., 2020; Yao et al., 2022).
@@
Transformer. In the past few years, transformer (Vaswani et al., 2017) has emerged as a dominant architecture for a wide range of applications, exhibiting their profound capacity in modeling different modalities: languages (Devlin et al., 2018; Liu et al., 2019; Touvron et al., 2023), vision (Dosovitskiy et al., 2020; Chen et al., 2021) and speech (Radford et al., 2023; Wang et al., 2023). Due to its ability to process multiple modalities, transformer-based architectures were widely adopted for image understanding using different types of supervision. For example, Vision Transformer (ViT) (Dosovitskiy et al., 2020) utilizes masked patches for pre-training and image-label supervision for fine-tuning. Notably, Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021), which was pre-trained on billion-scale image-text pairs, has demonstrated excellent zero-shot performance in image classification tasks. Despite obtaining impressive results on various downstream tasks, over-parameterization is still a crucial problem for transformer (Fan et al., 2019; Panahi et al., 2021) which potentially causes overfitting (Li et al., 2023) and unintended biases (Agarwal et al., 2021; Wang et al., 2021; Du et al., 2022; Zhang & Ré, 2022) more severely than small-scale architectures (Sagawa et al., 2020).
@@
Parameter-Efficient Fine-Tuning. While conventional training updates the entire network, prior work has shown that it is possible to achieve performance comparable with full fine-tuning by instead updating a small set of parameters (Lialin et al., 2023; He et al., 2023, 2021). Early work in this line of research includes adapter (Houlsby et al., 2019; Chen et al., 2022a; Sung et al., 2022), which adds a lightweight fully connected network between the layers of a frozen pre-trained model. Since then, there have been many other methods that update subsets of the network (Guo et al., 2021; Gheini et al., 2021; Zaken et al., 2022) or reparametrize the weights of the network using low-rank decompositions (Aghajanyan et al., 2020; Hu et al., 2021; Edalati et al., 2022). Another appealing PEFT strategy is prompt tuning, which prepends a few trainable tokens before the input sequence or hidden state to encode task-specific knowledge to the pre-trained model (Lester et al., 2021; Li & Liang, 2021; Zhu et al., 2023b; Huang et al., 2023).
Gradient-based multi-task learning. Prior work often searches for an update direction that benefits all tasks (Yu et al., 2020; Liu et al., 2021a; Phan et al., 2022a) or controls the trade-offs between per-task performance (Lin et al., 2019; Mahapatra & Rajan, 2020; Phan et al., 2022b). Therefore, casting the problem of learning from multiple domains into a multi-task problem (Liang et al., 2021; Kim et al., 2023) allows us to apply results from multi-objective optimization theory. Building on these insights, we propose a method that prevents the dominance of any single domain.
@@ 3.1 Problem formulation
Following previous methods (Sagawa et al., 2020), we adopt two metrics: worst group accuracy, indicating the minimum test accuracy across all groups $g\in\mathcal{G}$, and average accuracy, which represents the weighted average test accuracy with the weights corresponding to the relative proportions of each group in the training data. It is worth noting that commonly used datasets for group robustness often exhibit skewed training set distributions, the weighted average score is thus dominated by the performance of the model on groups that include large numbers of training data. Therefore, we propose to evaluate the model performance by additionally reporting the mean (unweighted average) accuracy score across different groups.
@@ 3.3 Multi-objective Optimization
In general, an optimal solution for this objective vector function does not exist since each of the individual function $f_{i}(x)$ does not necessarily guarantee to have the same minimum solution. Alternatively, we expect to obtain a solution, from which we cannot improve any specific objective without hurting another. According to the above formulation, the solution (Zitzler & Thiele, 1999) of the multi-objective minimization problem is formally defined as follows:
Definition 3.1.
@@
Gradient-based MOO. Désidéri (2012) shows that the descent direction can be found in the convex hull composed of gradients corresponding to different objectives $\{\boldsymbol{h}^{i}:=\nabla f_{i}\}_{i=1}^{m}$:
@@ 4 Proposed method
In GroupDRO, the updating formula solely focuses on the worst-performing group, which is susceptible to challenges when groups exhibit varying levels of loss magnitude, noise, and transfer characteristics. For example, among the high-loss group, there might exist a few outliers that have large input shifts or even wrong labels, which can raise the training loss of the whole group and prevent the model from fitting other groups.
@@ 4.1 Balancing group distributional robustness
Figure 2: Overview of our method on the Waterbirds dataset. Our main objective is to not only improve model performance across groups by optimizing their loss functions $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$, but also maximize the entropy over this loss distribution.
@@
where $\ell_{k}(\theta)=\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]$ and $w\in\Delta_{{}_{K}}$.
@@
However, MGDA is often biased toward objectives with smaller gradient magnitudes (Liu et al., 2020), and lacks controllable property (i.e. could not adjust the group that we want to focus more or less at each update), which can be utilized to balance the loss between groups. Hence, we introduce a new objective that steers the updating direction to obtain our desired balanced magnitude among groups. Assuming that the entropy $H(p)$ measures the diversity of a continuous distribution $p$, this behavior could be achieved via maximization of the entropy with respect to the distribution of loss functions:
$\mathcal{L}_{ent}(\theta)$ $=H\big{(}\text{softmax}\big{(}\vec{\mathcal{L}}(\theta))\big{)}$
$=H\Big{(}\text{softmax}\big{(}\big{\{}\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]\big{\}}_{k=1}^{K}\big{)}\Big{)}$ (5)
Theorem 4.1.
@@
The proof of Theorem 4.1 is deterred to Appendix A. Theorem 4.1 provides us with the updating rule that increases the entropy term in Equation (5) the most. To balance the learning across groups, we seek the coefficient $w$ that maximizes the cosine similarity with $d_{ent}$ and does not conflict against the gradient vector of each group. Denotes $w^{\prime}_{i}=p_{i}\log(p_{i})-p_{i}\sum_{j=1}^{K}\log(p_{j})p_{j}$, this is equivalent to optimizing:
@@
While the balance between loss magnitude is not achieved, we strictly follow the updating direction that decreases training errors of high-loss groups, but slightly relax constraints in the optimization problem (6) to allow bounded increases among groups that already have small loss magnitudes.
@@
In summary, the key difference between our proposed method and GroupDRO is that we solve a small $K$-dimensional optimization problem to find the updating direction that not only focuses on the worst group but also improves the performance in other groups and balances the loss magnitude between them at the same time. In practical settings, the group number $K$ is often a small number (e.g. $K=4$), thus could be efficiently solved using standard linear programming libraries (Diamond & Boyd, 2016; Andersen et al., 2020) without introducing significant computational overhead.
@@ 4.2 Controllable prompt tuning for balancing group distributional robustness
To achieve the controllable property, we apply the optimization procedure in section 4.1 on $\ell_{k}(\theta)={c}_{k}\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]$ where $\mathbf{c}\in\mathbb{R}^{K}$ is a pre-defined vector that is used to adjust the magnitude of group losses. Intuitively, the principle behind our proposed objective is to enforce the loss vector $\vec{\mathcal{L}}(\theta)$ to be inversely proportional to $\mathbf{c}$, from which we can adjust the group loss magnitude, and thus control the trade-off between them by varying $\mathbf{c}$. For example, we can put more weight on penalizing loss terms of some specific groups to accelerate the learning progress on those groups, or conversely slow down the learning to avoid overfitting on oversampled groups (Deng et al., 2023).
@@
Overview of our proposed method is depicted in Figure 2, given an input image belonging to one of a group $\{g^{1},g^{2},\dots,g^{K}\}$ and a set of classes, we solve the optimization problem (4.1) to find the reweighting coefficient $w$ that maximizes the entropy $\mathcal{L}_{ent}(\theta)$ while reasonably decreases group losses. Please note that while Figure 2 illustrates our proposed method on the CLIP model on the image classification task, it is applicable to other transformer-based models and task types. In our experiment, we exploit both image-end prompts and language-end prompts tuning on frozen ViT and CLIP backbones, respectively.
Prompt tuning for Vision Transformer: Regarding the prompt design for a ViT of $N$ layers, we use the set of continuous learnable tokens $\{\mathbf{P}_{n}\}_{n=0}^{N-1}$. Each $\mathbf{P}_{n}\in\mathbb{R}^{L\times D}$, where $L$, $D\in\mathbb{N}$ indicate the prompt length and latent space dimension, respectively. Subsequently, the input at the $(n+1)$-th layer follows the format of $\left[x_{n},\mathbf{P}_{n},\mathbf{E}_{n}\right]$, where $\mathbf{x}_{n},\mathbf{E}_{n}\in\mathbb{R}^{D}$ represent the [CLASS] token and image patch embeddings after $n$-th layer while [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operator. Thus, the number of trainable parameters for this image-end prompt set is $N\times L\times D$.
Prompt tuning for CLIP: Being trained on an enormous amount of image-text pairs, CLIP offers an expressive representation for both language and image inputs by using two separate encoders, one for each modality, as shown in Figure 2. Since the image encoder can be either ResNet (He et al., 2016), ConvNeXt (Liu et al., 2022) or ViT (Dosovitskiy et al., 2020), we conduct prompt tuning on the transformer text encoder for consistency. In short, we employ a learnable context prompt of length $L$: $\mathbf{P}\in\mathbb{R}^{L\times D}$, to replace the hand-crafted prompt used by the original CLIP (i.e., “this is a photo of a ”). Based on this design, the input for the text encoder of each class has the format of $\left[x,\mathbf{P}\right]$, where $x$ is the [CLASS] embedding of the corresponding class name and has the same dimension of $D$. These learnable tokens $\mathbf{P}$ are better at capturing domains-specific knowledge than those artificial prompt (Zhou et al., 2022; Zhu et al., 2023a) and can be learned to de-bias the original CLIP model while introducing a few trainable parameters ($L\times D$).